OpenStudy (anonymous):

square root of 398 solving calculus way

OpenStudy (anonymous):

go to calculatersoup.com search square roots

OpenStudy (anonymous):

what does it mean "calculus way"? maybe finding the linearization of \[f(x)=\sqrt{400+x}\] at x = 0?

OpenStudy (anonymous):

if so find \[f(0)=20\] \[f'(x)=\frac{1}{2\sqrt{400+x}}\] \[f'(0)=\frac{1}{40}\] and so the line is \[y=\frac{1}{40}x+20\]

OpenStudy (anonymous):

then for 389 put \[400+x=389\] \[x=-2\] and so your linear approximation is \[y=\frac{-2}{40}+20\] or \[-.05+20=19.95\]